Integration - Velocity/Acceleration Problem (with tomatoes)....

You are looking out your window in a skyscraper, and again your window is at a height of 450 feet above the ground. This time, however, you know that the tomato wasdropped (so v0 = 0), but you did not see it dropped so you do not know when it was dropped. You measure that it takes exactly 2 seconds from the time the tomato passesyour window until it hits the ground. From what height did it fall? Give your answer as a decimal approximation with units.
Answer #1

Distance travelled in 2 sec = 450

450 = u + 1/2 at2

t=2 a= g= 32 ft/sec2

386 =u

Let height of point of fall above window = h

u2=2gh

h= 2328.06 ft

Total height = h + 450 = 2778.06 ft

answered by: namuun